# Improved High-probability Convergence Guarantees of Decentralized SGD

### v2

@@ Abstract
Convergence in high-probability (HP) has been receiving increasing interest, due to its attractive properties, such as exponentially decaying tail bounds and strong guarantees for each individual run of an algorithm. While HP guarantees are extensively studied in centralized settings, much less is understood in the decentralized, networked setup. Existing HP studies in decentralized settings impose strong assumptions, like uniformly bounded gradients, or asymptotically vanishing noise, resulting in a significant gap between assumptions used to establish convergence in the HP and the mean-squared error (MSE) sense, even for vanilla Decentralized Stochastic Gradient Descent (DSGD) algorithm. This is contrary to centralized settings, where it is known that SGD converges in HP under the same conditions on the cost function as needed to guarantee MSE convergence. Motivated by this observation, we revisit HP guarantees for DSGD in the presence of light-tailed noise. We show that DSGD converges in HP under the same conditions on the cost as in the MSE sense, removing uniformly bounded gradients and other restrictive assumptions, while simultaneously achieving order-optimal rates for both non-convex and strongly convex costs. Moreover, our improved analysis yields linear speed-up in the number of users, demonstrating that DSGD maintains strong performance in the HP sense and matches existing MSE guarantees. Our improved results stem from a careful analysis of the MGF of quantities of interest (norm-squared of gradient or optimality gap) and the MGF of the consensus gap between users’ models. To achieve linear speed-up, we provide a novel result on the variance-reduction effect of decentralized methods in the HP sense and more fine-grained bounds on the MGF for strongly convex costs, which are both of independent interest.
@@ 1 Introduction
Modern large-scale machine learning applications and the abundance of data necessitate alternatives to the centralized computation framework, giving rise to distributed learning, a paradigm where multiple users collaborate to jointly train a model, e.g., [sayed-networks, pmlr-v54-mcmahan17a, vlaski_et_al]. The features of distributed learning, such as storing the data locally and only exchanging smaller updates, like (quantized) model parameters or gradients, and the lack of a single point of failure, further make it an attractive paradigm from a privacy and security perspective [bonawitz2017practical, yu-secure]. Many applications, such as federated training of models on mobile devices [konecny2016federated], controlling and coordinating robot swarms [bullo2009distributed], or distributed control and power grids [dist-learn-noncvx], all rely on distributed computation. From a communication/model exchange perspective, distributed frameworks can be client-server (i.e., federated) or decentralized (i.e., networked), with the main difference being that in the client-server setup users communicate with a server, while in decentralized settings users communicate directly with each other. [^1] While the client-server setup is often characterized by performing local updates and periodic communication [stichlocal], as well as partial user participation [pmlr-v151-jee-cho22a], we note that these features can also be incorporated in the decentralized setup, by modifying the methods to perform multiple local updates before communicating [pmlr-v119-koloskova20a] and considering a dynamic network with users that are occasionally idle [Cannelli2020]. However, this is not the focus of our current work. Noting that, from the model update perspective, the client-server setup is equivalent to the decentralized setup with a fully connected communication network, we focus on the more general, decentralized setup.
The study of convergence guarantees of decentralized optimization algorithms has a long history, e.g., [nedic-subgrad, jakovetic-fast, shi-extra, scutari-next, kun-dgd, jakovetic-unification, swenson-dsgd], with most works focusing on MSE convergence, e.g., [conv-rates-dsgd-jakovetic, Pu2021, vlaski-dsgd-nonconv-1, vlaski-dsgd-nonconv-2, pmlr-v119-koloskova20a, ran-vr, ran-improved, wang-cooperative], see also [adaptive-learning-ali] for an extensive treatment of the topic. Another type of convergence guarantees, namely convergence in HP, has garnered increasing attention recently. In particular, for a non-negative stochastic process $\{X^{t}\}_{t\in\mathbb{N}}$, the goal of HP convergence guarantees is to establish, for all $t\in\mathbb{N}$ and any $\epsilon>0$, that
@@
where $\gamma_{1},\gamma_{2},C>0$ are some positive constants. If $\{X^{t}\}_{t\in\mathbb{N}}$ is a measure of performance of an algorithm, e.g., $X^{t}=\frac{1}{t}\sum_{k\in[t]}\|\nabla f(x^{k})\|^{2}$, where $\{x^{t}\}_{t\in\mathbb{N}}$ is a sequence generated by the algorithm and $f$ is a non-convex cost, the relation (1) provides strong guarantees with respect to a single run of the algorithm. This is particularly important in modern applications like LLMs, where it is often intractable to perform multiple runs. Numerous works study HP guarantees of SGD-type methods in centralized settings, under both light-tailed [nemirovski2009robust, ghadimi2013stochastic, li-orabona, pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high] and heavy-tailed noise [nguyen2023improved, liu2023breaking, hubler2024normalization, kornilov2025sign, armacki2023high]. Comparatively, there have been few studies on HP convergence of decentralized methods.
@@ 1.1 Literature Review
We now review the literature, focusing on works studying centralized HP convergence guarantees and decentralized MSE and HP convergence guarantees.
Centralized HP Convergence.
@@
The authors in [nemirovski2009robust] show optimal convergence rates of SGD for convex costs under light-tailed (i.e., sub-Gaussian) stochastic gradients (see assumption (A4) ahead for a formal definition of sub-Gaussianity), while the work [ghadimi2013stochastic] shows the same for non-convex costs. The authors in [li-orabona, li2020high] establish high-probability guarantees of momentum SGD and a variant of AdaGrad for non-convex costs, respectively. The works [pmlr-v99-harvey19a] and [pmlr-v206-bajovic23a] respectively provide optimal HP convergence rates for the last iterate of SGD for non-smooth and smooth strongly convex costs, while in [liu2024revisiting] the authors establish unified HP convergence guarantees for smooth and non-smooth convex and strongly convex costs. The work [liu2023high] generalizes the previous studies for non-convex and convex costs, providing unified guarantees for several algorithms, including SGD and AdaGrad for smooth and non-smooth costs. The authors in [madden2020high] study HP convergence of SGD under sub-Weibull noise. Another line of work studies HP convergence under heavy-tailed noise, [^2] Since heavy-tailed noise is not the focus of our work, we point the reader to [heavy-tail-book, nguyen2023improved, armacki2023high] for various conditions used when studying heavy-tailed noise. e.g., [sadiev2023highprobability, nguyen2023improved, liu2023breaking, gorbunov2023breaking, hubler2024normalization, kornilov2025sign, armacki2024_ldp+mse, armacki2025optimal, armacki2023high], where it is necessary to introduce algorithmic modifications, e.g., clipping, normalization, or sign, to ensure concentration of the form in (1). Crucially, convergence in the HP sense is achieved under the same conditions on the cost function as in the MSE sense, for both light-tailed and heavy-tailed noise.
Decentralized MSE Convergence.
@@
MSE guarantees are typically studied under a variety of bounds on the second noise moment. [^3] See, e.g., [khaled2022better] for a good overview of the various conditions used in the literature. Following one of these settings, the work [conv-rates-dsgd-jakovetic] shows that DSGD converges at an optimal rate for strongly convex costs, while in [Pu2021] the authors show that DSGD with a fixed step-size and the gradient tracking (GT) mechanism, e.g., [scutari-next, nedic-tracking, qu-harnessing], converges to a neighbourhood of the optimal solution. The works [vlaski-dsgd-nonconv-1, vlaski-dsgd-nonconv-2, vlaski-dsgd-nonconv-3] study MSE guarantees of DSGD for non-convex costs and show that it escapes saddle points with high probability. In [wang-cooperative] the authors propose a general framework dubbed cooperative SGD, showing optimal rates for non-convex costs. The work [koloskova2023grad_clip] provides unified guarantees for DSGD with local updates and changing network topology, with optimal rates and linear speed-up in the number of users for non-convex and (strongly) convex costs. In [ran-vr] the authors show DSGD with GT and variance reduction converges at a linear rate for strongly convex costs, while in [ran-improved] optimal rates of DSGD with GT and linear speed-up in the number of users are established for non-convex costs and costs satisfying the Polyak-Łojasiewicz (PL) condition [karimi2016linear]. Finally, in [yu2025decentralized] the authors show that DSGD with GT, local momentum and gradient normalization achieves optimal rates under heavy-tailed noise. It is also worth mentioning a rich line of works studying MSE guarantees in decentralized settings for problems such as estimation and detection, multi-objective and multitask optimization, see [consensus+innovation1, chen-decent-pareto, nassif-multitask] and references therein.
Decentralized HP Convergence.
@@
Compared to MSE guarantees, there is a significantly smaller body of work on HP convergence guarantees of decentralized algorithms. In particular, the work [dsgd-high-prob] studies HP convergence of DSGD for both non-convex and PŁ costs under light-tailed noise, requiring uniformly bounded gradients and asymptotically vanishing noise for PŁ costs (see discussion after Theorem 2 ahead), showing optimal rates in both cases. In [dsgd-online-noncooperative-games, dsgd-online-stochastic] the authors study HP convergence of a decentralized mirror descent algorithm under sub-Gaussian noise and bounded gradients, for online noncooperative games and dynamic regrets, respectively. Works [clipped-dual-avg, clipped-dsgd1, clipped-dsgd2] study HP convergence of decentralized algorithms with clipping under heavy-tailed noise. It is also worth mentioning [bajovic-detection-ld, bajovic-detection-ld2, matta-diffusion-1, matta-difussion-2, bajovic-inference-ld], where the authors study asymptotic large deviation guarantees for decentralized problems like detection and inference. A common thread for all these works is the need for uniformly bounded gradients, or algorithmic modifications like gradient clipping, which ensure that the gradient stays bounded. This observation raises the question of whether decentralized algorithms can converge in the HP sense without ensuring that the gradient is uniformly bounded, creating a gap between convergence in the HP and MSE sense, where bounded gradients are not required. This is further contrasted by the centralized setting, where it is known that SGD-type algorithms require the same conditions on the cost function to guarantee both HP and MSE convergence. Moreover, none of the existing HP convergence bounds in decentralized settings achieve linear speed-up in the number of users, which is a known feature of decentralized algorithms, established in the MSE sense in, e.g., [pmlr-v119-koloskova20a, ran-improved, unified-refined].
@@ 1.2 Contributions
Motivated by the observed gap between the existing literature on HP and MSE convergence in decentralized settings, we revisit the HP convergence guarantees in decentralized stochastic optimization, by studying a variant of vanilla DSGD in the presence of light-tailed noise. We establish that DSGD converges in HP under the same conditions on the cost as required in the MSE sense, for both non-convex and strongly convex costs, removing the uniformly bounded gradient condition. Next, we show that DSGD achieves optimal rates and linear speed-up in the number of users, for both non-convex and strongly convex costs, further improving on existing works and closing the gap between HP and MSE guarantees in decentralized settings. Our results are established by carefully bounding the MGF of the quantity of interest (i.e., average norm-squared of the gradient for non-convex and optimality gap of the last iterate for strongly convex costs) and the MGF of the consensus gap. Compared to [dsgd-high-prob], the work closest to ours, we provide several improvements. In particular, we remove the uniformly bounded gradient requirement, as well as the vanishing noise requirement imposed for PŁ costs (see the discussion after Theorem 2 ahead), while achieving linear speed-up in the number of users. Compared to works studying MSE guarantees for DSGD, e.g., [pmlr-v119-koloskova20a, conv-rates-dsgd-jakovetic], our results require directly working with the MGF and carefully balancing between the MGF of the quantity of interest and that of the consensus gap. This is particularly challenging for strongly convex costs, where we want to show an “almost decreasing” property of the MGF, in order to get improved rates on the last iterate (see Lemma 5 and the related discussion ahead). Compared to centralized SGD, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high, liu2024revisiting], the main challenge lies in the additional consensus gap and simultaneously dealing with its MGF and that of the quantity of interest. All of these challenges are resolved by introducing several novelties, outlined next.
@@
Toward establishing our improved HP convergence guarantees, we face several technical challenges, requiring novel results. In particular, to remove uniformly bounded gradients, we provide Lemma 3 and use the “offset trick” for non-convex costs (see the proof sketch of Theorem 1 in Section 4 for details), while for strongly convex costs we establish a novel bound on the MGF of the consensus gap, in the form of Lemma 4. To establish linear speed-up in the number of users, we show that the variance reduction benefit of decentralized learning is maintained in the HP sense, in the form of Lemma 2, which is of independent interest when studying HP guarantees of decentralized algorithms. Further, to guarantee linear speed-up for strongly convex costs, we require a more fine-grained bound on the MGF, provided in the form of Lemma 5, which allows for incorporating higher-order terms, generalizing the bounds from [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2024revisiting] and is of independent interest, even in centralized settings.
@@
The rest of the paper is organized as follows. Section 2 formally introduces the problem and the DSGD method, Section 3 outlines the main results, Section 4 provides proof sketches and discussion, while Section 5 concludes the paper. Appendix contains results omitted from the main body. The remainder of this section introduces the notation.
@@
We use $\mathbb{N}$, $\mathbb{R}$ and $\mathbb{R}^{d}$ to denote positive integers, real numbers and $d$-dimensional vectors. For $m\in\mathbb{N}$, we use $m]=\{1,\ldots,m\}$ to denote positive integers up to and including $m$. The notation $\langle\cdot,\cdot\rangle$ stands for the Euclidean inner product, while $\|\cdot\|$ is used for both the induced vector and matrix $\ell_{2}$ norms. We use subscripts to denote users and superscripts to denote the iteration counter, e.g., $x^{t}_{i}$ refers to the model of user $i$ in iteration $t$. We use the “big O” notation $\mathcal{O}(\cdot)$ to hide global constants, unless stated otherwise.
2 Problem Setup and Proposed Method

### v3

# High-probability Convergence Guarantees of Decentralized SGD

@@ Abstract
Convergence in high-probability (HP) has attracted increasing interest, due to implying exponentially decaying tail bounds and strong guarantees for individual runs of an algorithm. While many works study HP guarantees in centralized settings, much less is understood in the decentralized setup, where existing works require strong assumptions, like uniformly bounded gradients, or asymptotically vanishing noise. This results in a significant gap between the assumptions used to establish convergence in the HP and the mean-squared error (MSE) sense, and is also contrary to centralized settings, where it is known that SGD converges in HP under the same conditions on the cost function as needed for MSE convergence. Motivated by these observations, we study the HP convergence of Decentralized SGD (DSGD) in the presence of light-tailed noise, providing several strong results. First, we show that DSGD converges in HP under the same conditions on the cost as in the MSE sense, removing the restrictive assumptions used in prior works. Second, our sharp analysis yields order-optimal rates for both non-convex and strongly convex costs. Third, we establish a linear speed-up in the number of users, leading to matching, or strictly better transient times than those obtained from MSE results, further underlining the tightness of our analysis. To the best of our knowledge, this is the first work that shows DSGD achieves a linear speed-up in the HP sense. Our relaxed assumptions and sharp rates stem from several technical results of independent interest, including a result on the variance-reduction effect of decentralized methods in the HP sense, as well as a novel bound on the MGF of strongly convex costs, which is of interest even in centralized settings. Finally, we provide experiments that validate our theory.
@@ 1 Introduction
Modern large-scale machine learning applications and the abundance of data necessitate alternatives to the centralized computation framework, giving rise to distributed learning, a paradigm where multiple users collaborate to jointly train a model, e.g., [63, 49, 68]. The features of distributed learning, such as storing the data locally and only exchanging smaller updates, like (quantized) model parameters or gradients, and the lack of a single point of failure, further make it an attractive paradigm from a privacy and security perspective [10, 80]. Many applications, such as federated training of models on mobile devices [34], controlling and coordinating robot swarms [11], or distributed control and power grids [14], all rely on distributed computation. From a communication perspective, distributed frameworks can be client-server (i.e., federated) or decentralized (i.e., networked), with the main difference being that in the client-server setup users communicate with a central server, while in decentralized settings users communicate directly with each other. [^1] While the client-server setup is often characterized by performing local updates and periodic communication [65], as well as partial user participation [17], we note that these features can also be incorporated in the decentralized setup, by modifying the methods to perform multiple local updates before communicating [33] and considering a dynamic network with users that are occasionally idle [12]. However, this is not the focus of our current work. Noting that, from the model update perspective, the client-server setup is equivalent to the decentralized setup with a fully connected communication network, we focus on the more general, decentralized setup.
The study of convergence guarantees of decentralized optimization algorithms has a long history, e.g., [53, 25, 64, 43, 81, 26, 66], with most works focusing on MSE convergence, e.g., [24, 57, 70, 69, 33, 74, 75, 72], see also [62] for an extensive treatment of the topic. Another type of convergence guarantees, namely convergence in HP, has garnered increasing attention recently. In particular, for a non-negative stochastic process $\{X^{t}\}_{t\in\mathbb{N}}$, the goal of HP convergence is to establish that, for all $t\in\mathbb{N}$ and any $\epsilon>0$
@@
where $\gamma_{1},\gamma_{2},C>0$ are constants. If $\{X^{t}\}_{t\in\mathbb{N}}$ is a measure of performance, e.g., $X^{t}=\frac{1}{t}\sum_{k\in[t]}\|\nabla f(x^{k})\|^{2}$, where $\{x^{t}\}_{t\in\mathbb{N}}$ is a sequence generated by some algorithm and $f$ is a non-convex cost, the relation (1) provides strong guarantees with respect to a single run of that algorithm. This is particularly important in modern applications like LLMs, where it is often intractable to perform multiple runs. Numerous works study HP guarantees of SGD-type methods in centralized settings, under both light-tailed [54, 20, 36, 21, 5, 39] and heavy-tailed noise [56, 40, 23, 35, 4]. Comparatively, there have been very few studies of HP convergence of decentralized methods.
@@ 1.1 Literature Review
We now review the literature, focusing on centralized HP results and decentralized MSE and HP convergence results.
@@
Centralized HP Convergence. [54] show optimal convergence rates of SGD for convex costs under light-tailed stochastic gradients (see assumption (A4) ahead for a formal definition of light-tailed noise), while [20] show the same for non-convex costs. [36, 37] establish high-probability guarantees of momentum SGD and a variant of AdaGrad for non-convex costs, respectively. [21] and [5] respectively provide optimal HP convergence rates for the last iterate of SGD for non-smooth and smooth strongly convex costs, while [41] establish unified HP convergence guarantees for smooth and non-smooth convex and strongly convex costs. [39] generalize the previous works on non-convex and convex costs, providing unified guarantees for several algorithms, including SGD and AdaGrad for smooth and non-smooth costs. [46] study HP convergence of SGD under sub-Weibull noise. Another line of work studies HP convergence under heavy-tailed noise, [^2] Since heavy-tailed noise is not the focus of our work, we point the reader to [50, 56, 4] for various conditions used to study heavy-tailed noise. e.g., [61, 56, 40, 58, 23, 35, 3, 2, 4], where it is necessary to introduce algorithmic modifications, e.g., clipping, normalization, or sign, to ensure concentration of the form in (1). Crucially, convergence in the HP sense is achieved under the same conditions on the cost function as in the MSE sense, for both light-tailed and heavy-tailed noise.
@@
Decentralized MSE Convergence. MSE guarantees are typically studied under a variety of bounds on the second noise moment. [^3] See, e.g., [31] for a good overview of the various conditions used in the literature. Following one of these settings, [24] show that DSGD converges at an optimal rate for strongly convex costs, while [57] show that DSGD with a fixed step-size and the gradient tracking (GT) mechanism, e.g., [43, 52, 60], converges to a neighbourhood of the optimal solution. [70, 69, 71] study MSE guarantees of DSGD for non-convex costs and show that it escapes saddle points with high probability. [72] propose a general framework dubbed cooperative SGD, showing optimal rates for non-convex costs. [32] provide unified guarantees for DSGD with local updates and changing network topology, with optimal rates and linear speed-up in the number of users for non-convex and (strongly) convex costs. [74] show DSGD with GT and variance reduction converges at a linear rate for strongly convex costs, while [75] establish optimal rates of DSGD with GT and linear speed-up in the number of users for non-convex costs and costs satisfying the Polyak-Łojasiewicz (PL) condition, e.g., [30]. [79] show that DSGD with GT, momentum and normalization achieves optimal rates under heavy-tailed noise. It is worth mentioning a rich line of works studying MSE guarantees for decentralized problems such as estimation, detection, multi-objective and multitask optimization, see, e.g., [28, 16, 51] and references therein.
@@
Decentralized HP Convergence. Compared to MSE guarantees, there is a significantly smaller body of work on HP convergence guarantees of decentralized algorithms. In particular, [44] study HP convergence of DSGD for general non-convex and PŁ costs under light-tailed noise, requiring uniformly bounded gradients, as well as asymptotically vanishing noise for PŁ costs (see the discussion after Theorem 2 for details), while showing optimal rates in both cases. [45, 76] study HP convergence of a decentralized mirror descent algorithm under light-tailed noise, for online noncooperative games and dynamic regrets, respectively, requiring bounded gradients and compact domains. [59, 77] study HP convergence of decentralized algorithms with clipping, for convex costs under heavy-tailed noise, while [78] also consider general non-convex costs. It is worth mentioning works like [7, 6, 47, 48, 8], where the authors study asymptotic large deviation guarantees for decentralized problems such as detection and inference. A common thread for all these works is the need for uniformly bounded gradients, or algorithmic modifications like gradient clipping, which ensure that the gradients stay bounded, as well as the fact that none of the existing HP studies achieve linear speed-up in the number of users, which is a known feature of decentralized algorithms in the MSE sense in, e.g., [33, 75, 1]. These observations raise the following important questions: (i) Can decentralized algorithms converge in the HP sense under the same assumptions on the cost as used for MSE results, where bounded gradients are not required? (ii) Is it possible to achieve a linear speed-up in the number of users in the HP sense, which is a well-known feature of decentralized algorithms? This situation is further contrasted by the centralized setting, where it is known that SGD-type algorithms require the same conditions on the cost function to guarantee both HP and MSE convergence.
@@ 1.2 Contributions
Motivated by the observed gaps between the existing literature on HP and MSE convergence in decentralized settings, we revisit the HP convergence guarantees in decentralized optimization, by studying a variant of vanilla DSGD in the presence of light-tailed noise, establishing several strong results in the process. In particular, we show that DSGD converges in HP under the same conditions on the cost as required in the MSE sense, for both non-convex and strongly convex costs, removing strong conditions like uniformly bounded gradients and (asymptotically) vanishing noise. Next, we show that DSGD achieves optimal rates and linear speed-up in the number of users, for both non-convex and strongly convex costs, further improving on existing works and closing the gap between HP and MSE guarantees in decentralized settings. Our results are established by carefully bounding the MGF of the quantity of interest (i.e., average norm-squared of the gradient for non-convex and optimality gap of the last iterate for strongly convex costs) and the MGF of the consensus gap. Compared to [44], the work closest to ours, we provide several improvements. In particular, we remove the uniformly bounded gradient requirement, as well as the vanishing noise requirement imposed for PŁ costs (see the discussion after Theorem 2 ahead), while achieving linear speed-up in the number of users. Compared to works studying MSE guarantees for DSGD, e.g., [33, 24], our results require directly working with the MGF and carefully balancing between the MGF of the quantity of interest and that of the consensus gap. This is particularly challenging for strongly convex costs, where we want to show an “almost decreasing” property of the MGF, in order to get improved rates on the last iterate (see Lemma 5 and the related discussion ahead). Compared to centralized SGD, e.g., [21, 5, 39, 41], the main challenge lies in the additional consensus gap and simultaneously dealing with its MGF and that of the quantity of interest. These challenges are resolved by introducing several novelties, outlined next.
@@
Toward establishing our improved HP convergence guarantees, we face several technical challenges, requiring novel results. In particular, to remove uniformly bounded gradients, we provide Lemma 3 and use the “offset trick” for non-convex costs (see the proof sketch of Theorem 1 in Section 4 for details), while for strongly convex costs we establish a novel bound on the MGF of the consensus gap, removing the need for bounded gradients and bounded heterogeneity, in the form of Lemma 4. To establish linear speed-up in the number of users, we show that the variance reduction benefit of decentralized learning is maintained in the HP sense, in the form of Lemma 2, which is of independent interest when studying HP guarantees of decentralized algorithms. Further, we establish a novel result on the MGF of the optimality gap for strongly convex costs (more broadly, “almost decreasing” processes, see Section 3.3 ahead), in the form of Lemma 5, which is essential in ensuring linear speed-up, by providing a more fine-grained bound on the MGF compared to existing centralized results [21, 5, 41] and is of independent interest, even in centralized settings.
@@
The rest of the paper is organized as follows. Section 2 outlines the problem and DSGD method, Section 3 presents the main results, Section 4 provides proof sketches and discussions, while Section 5 concludes the paper. Appendix contains results omitted from the main body. The remainder of this section introduces the notation.
@@
We use $\mathbb{N}$, $\mathbb{R}$ and $\mathbb{R}^{d}$ to denote positive integers, real numbers and $d$-dimensional vectors. For $m\in\mathbb{N}$, we use $m]=\{1,\ldots,m\}$ to denote positive integers up to and including $m$. The notation $\langle\cdot,\cdot\rangle$ stands for the Euclidean inner product, while $\|\cdot\|$ is used for both vector and matrix induced norms. We use subscripts to denote users and superscripts to denote the iteration counter, e.g., $x^{t}_{i}$ refers to the model of user $i$ in iteration $t$. The “big O” notation $\mathcal{O}(\cdot)$ hides only global constants, unless stated otherwise.
2 Problem Setup and DSGD